OpenStudy (anonymous):

I BET NOONE CAN ANSWER THIS: One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (saifoo.khan):

And wht if someone solves it?

OpenStudy (chaise):

Let's assume that HM = Human and HR = Horse HM + HR = 74 2HM + 4HR = 196 (2HM + 4HR) - (2 HM + 2HR) = 196 - 148 2HR = 48 HR = 24 HM + (24) = 74 HM = 74 - 24 HM = 50 So, the solution is 24 horses and 50 humans.

OpenStudy (saifoo.khan):

Chay-C won the bet.

OpenStudy (anonymous):

WOW how did you type that out so fast chaise..........

OpenStudy (anonymous):

chaise got it

OpenStudy (anonymous):

copy and paste

OpenStudy (saifoo.khan):

LOL^

OpenStudy (anonymous):

HAX I SAY HAX

OpenStudy (chaise):

Haha. This could easily been solved manually - but I figured you got the question from the internet - so the solution would be in the internet also. Quick Google search found it ;)

OpenStudy (anonymous):

ok wat about: y = log x If y = 10, then what is x?

OpenStudy (chaise):

x would be equal to 1. Assuming log is base 10, ofcourse. Everyone knows log of 10 is 1.

OpenStudy (saifoo.khan):

stop copying frmo Oracle.

OpenStudy (saifoo.khan):

from*

OpenStudy (anonymous):

X=1